Suppose that a hypothetical element consists of a mixture of two isotopes. One isotope, having mass 44 amu, is present in 30.1% abundance, while the other isotope, having mass 46 amu, accounts for the other 69.9%. What should be the experimentally deter- mined atomic weight for this hypothetical el- ement?
Answer in units of amu.
1
Expert's answer
2017-01-30T14:57:12-0500
Ar(E) = Ar(Isotope1)w(Isotope1) + Ar(Isotope2)w(Isotope2) = 440.301 + 460.699 = 13.244 + 32.154 = 45.398 amu